# Characterizing Dark Bosons at Chiral Belle

### v1

@@ IV.2 Model comparison
Using the analysis described above, the right panel of Fig. 2 shows the upper bounds for a $100~{}\rm MeV$ boson in the $g_{A}$ vs $g_{V}$ plane assuming $20~{}{\rm ab}^{-1}$ of $P=0.7$ (blue) and $P=-0.7$ (orange) collisions which corresponds to $40~{}{\rm ab}^{-1}$ of unpolarized data (black). We also show the relationship between $g_{V}$ and $g_{A}$ in the three benchmark models discussed above by the dashed green lines.
To show the distinguishing power of a polarized analysis, we consider the scenario of a future discovery with $m_{V}=0.1~{}\rm GeV$. We take the couplings such that $40~{}{\rm ab}^{-1}$ of unpolarized data show a $5\sigma$ excess compared to the background-only hypothesis. This corresponds to $\epsilon_{\rm eff}=(1.25\pm 0.13)\times 10^{-4}$. We show the resulting ±$\pm 1\sigma$ favored region in the $g_{V}$-$g_{A}$ plane in Fig. 3 as the gray shaded circular regions for each of the three benchmark models we consider. Each of these models implies a different expected number of events in the two $20~{}{\rm ab}^{-1}$ of data for each $e^{-}$ polarization. We show the ±$\pm 1\sigma$ expected couplings that would be extracted from the two polarization measurements as blue ($P=0.7$) and orange ($P=-0.7$) shaded regions. We see that the overlap of these analyses occurs in regions of the $g_{A}$ vs. $g_{V}$ plane that correspond to the assumed underlying models (cf. right panel of Fig. 2).
@@
Because the dark photon and dark $Z$ cases both have approximately a single parity of coupling to the electron, $g_{A}\ll g_{V}$ and $g_{V}\ll g_{A}$ respectively, the regions of overlap are near the axes in the $g_{V}$-$g_{A}$ plane. In contrast, the $V_{R}$ case leads to a large difference in the number of signal events between the two polarizations and can be distinguished from the previous two models. This can also hold for more generic dark vectors, possibly in cases where kinetic and mass mixing are comparable in importance.
@@ V Summary and Outlook
Our results build on and complement the already world-leading capabilities of Belle II for this class of dark matter models. The experiment’s high luminosity, clean environment, and ability to trigger on mono-photon events make it an ideal laboratory for precision measurements and rare searches. With the addition of polarization of the electron beam, Chiral Belle can go significantly further—not just in extending our understanding of the Standard Model, but opening the possibility to discern the fundamental structure of new dark states.
@@
A straightforward extension of this work involves searches for heavier dark states that are produced off-shell at a $B$ factory. These scenarios are also well motivated, and there is overlap with $e^{+}e^{-}\to\gamma Z^{\ast}\to\gamma\nu\bar{\nu}$, which itself could present interesting physics. We explore this scenario in detail in a follow-up work [40]. Such searches can be complementary to direct [44, 78, *CMS:2024jyb] or indirect [80, *Kribs:2020vyk, *Thomas:2021lub, *Thomas:2022qhj, *Yan:2022npz] searches at higher energy colliders.

### v2

@@ IV.2 Model comparison
Using the analysis described above, the right panel of Fig. 2 shows the upper bounds for a $100~\rm MeV$ boson in the $g_{A}$ vs $g_{V}$ plane assuming $20~{\rm ab}^{-1}$ of $P=0.7$ (blue) and P=0.7P=-0.7 (orange) collisions which corresponds to $40~{\rm ab}^{-1}$ of unpolarized data (black). We also show the relationship between $g_{V}$ and $g_{A}$ in the three benchmark models discussed above by the dashed green lines.
To show the distinguishing power of a polarized analysis, we consider the scenario of a future discovery with $m_{V}=3.67~\rm GeV$. We take the couplings such that $40~{\rm ab}^{-1}$ of unpolarized data show a $5\sigma$ excess compared to the background-only hypothesis. This corresponds to $\epsilon_{\rm eff}=(3.96\pm 0.4)\times 10^{-4}$. We show the resulting $\pm 1\sigma$ favored region in the $g_{V}$-$g_{A}$ plane in Fig. 3 as the gray shaded circular regions for each of the three benchmark models we consider. Each of these models implies a different expected number of events in the two $20~{\rm ab}^{-1}$ of data for each $e^{-}$ polarization. We show the $\pm 1\sigma$ expected couplings that would be extracted from the two polarization measurements as blue ($P=0.7$) and orange (P=0.7P=-0.7) shaded regions. We see that the overlap of these analyses occurs in regions of the $g_{A}$ vs. $g_{V}$ plane that correspond to the assumed underlying models (cf. right panel of Fig. 2).
@@
Because the dark photon and dark $Z$ cases both have approximately a single parity of coupling to the electron, $g_{A}\ll g_{V}$ and $g_{V}\ll g_{A}$ respectively, the regions of overlap are near the axes in the $g_{V}$-$g_{A}$ plane. In contrast, the $Z_{R}$ case leads to a large difference in the number of signal events between the two polarizations and can be distinguished from the previous two models. This can also hold for more generic dark vectors, possibly in cases where kinetic and mass mixing are comparable in importance.
@@ V Summary and Outlook
Our results build on and complement the already world-leading capabilities of Belle II for this class of dark matter models. The experiment’s high luminosity, clean environment, and ability to trigger on mono-photon events make it an ideal laboratory for precision measurements and rare searches. With the addition of polarization of the electron beam, Chiral Belle can go significantly further, not just in extending our understanding of the Standard Model, but opening the possibility to discern the fundamental structure of new dark states.
@@
A straightforward extension of this work involves searches for heavier dark states that are produced off-shell at a $B$ factory. These scenarios are also well motivated, and there is overlap with $e^{+}e^{-}\to\gamma Z^{\ast}\to\gamma\nu\bar{\nu}$, which itself could present interesting physics. We explore this scenario in detail in a follow-up work [30]. Such searches can be complementary to direct [28, 1, *CMS:2024jyb] or indirect [42, *Kribs:2020vyk, *Thomas:2021lub, *Thomas:2022qhj, *Yan:2022npz] searches at higher energy colliders.